OpenStudy (anonymous):

(16x^3+4x^2-144)÷(4x-8)

OpenStudy (mertsj):

\[\frac{16x^3+4x^2-144}{4x-8}=\frac{4(4x^3+x^2-36)}{4(x-2)}=\frac{4x^3+x^2-36}{x-2}\]

OpenStudy (anonymous):

thank you!

OpenStudy (mertsj):

Are you supposed to finish the division using long division or synthetic division or were you only supposed to reduce the fraction to lowest terms?

OpenStudy (anonymous):

i was suppose to use long division

OpenStudy (mertsj):